25 cm3 of a solution containing T grams of K2CO3 in 1000 cm3 of solution were completely neutralized by 20.00 cm3 of a solution containing 0.025 mole of H2SO4 in 1 dm3 of solution

Calculate the :

(a) concentration of K2CO3 solution in moldm-3

(b) Value of T

(c) mass of K2SO4 formed

(d) volume of CO2 evolved at STP

(e) number of H2O molecules


1
Expert's answer
2022-03-27T10:29:24-0400

Volume of solution = "25 cm^3" , T gram of "K_2CO_3" ,Volume of solution = 1000 cm3


(a

(a) Volume of "H_2SO_4 =" 20 cm3 , Mole of "H_2SO_4=" 0.025

Reaction between sulphuric acid and potassium carbonate

"K_2CO_3+H_2SO_4 \longrightarrow K_2SO_4+CO_2+H_2O"

1 mole of potassium carbonate reacts with one mole of sulphuric acid so 0.025 mole of sulphuric react with 0.025 mole of potassium carbonate


Using molarity equation



"n_1M_1V_1= n_2M_2V_2"


"0.025\times M_1\times 25= 0.025\times 0.025 \times 20"


M1= 0.02 M

(b)mole of potassium carbonate = 0.025


Mass of potassium carbonate = T = "0.025 \times 138=3.45 g"


(c) Mass of K2SO4 = "174.26 \times 0.025 = 4.3565 g"


(d) Mole of carbon dioxide produced at STP=0.025


Volume of carbon dioxide = "0.025 \times 22.4 =0.56 L"


(e) mole of water molecules produced = 0.025


number of water molecules = "6.023 \times 10^{23}\times 0.025 = 1.5 \times 10^{22}"